OpenStudy (anonymous):

What is the equation for a circle with a center of (0, 2) and a radius of 11?

OpenStudy (jack1):

x^2 + y^2 = a^2 is the equation for a circle where a is the radius

OpenStudy (campbell_st):

the general form of the circle is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius you know the centre is (0, 2) and r = 11 substitute and then simplify